OpenStudy (anonymous):

Part A: Divide (8x4y3 + 4x3y2 - 2x2y - 12x2y4) by -2x2y. Show your work, and justify each step. (6 points) Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? (2 points) Part C: What is the DEGREE and classification of the polynomial you got in Part A? (2 points)

OpenStudy (anonymous):

divide each term separately for example \[\frac{8x^4y^2}{-2x^2y}=-4x^2y\]

OpenStudy (anonymous):

divide the number, subtract the exponent do it three more times

OpenStudy (anonymous):

4x^2y^2 - 2xy +6y^3

OpenStudy (anonymous):

be careful with your signs

OpenStudy (anonymous):

you are dividing by \( -2x^2y\) you have to change all the signs

OpenStudy (anonymous):

oh wait i am confused what is your final answer?

OpenStudy (anonymous):

\[4x^2y^2 - 2xy +6y^3\] is what you wrote, and it is close but not quite

OpenStudy (anonymous):

you are missing the minus sign on the first term also \[\frac{-2x^2y}{-2x^2y}=1\] not zero you dropped that one

OpenStudy (anonymous):

\[-4x^2y^2 - 2xy+1+6y^3\] would be more like it